# SOUS VIDE: Cooking Visual Drone Navigation Policies in a Gaussian Splatting Vacuum

### v1

@@ Abstract
We propose a new simulator, training approach, and policy architecture, collectively called SOUS VIDE, for end-to-end visual drone navigation. Our trained policies exhibit zero-shot sim-to-real transfer with robust real-world performance using only on-board perception and computation. Our simulator, called FiGS, couples a computationally simple drone dynamics model with a high visual fidelity Gaussian Splatting scene reconstruction. FiGS can quickly simulate drone flights producing photorealistic images at up to 130 fps. We use FiGS to collect 100k-300k observation-action pairs from an expert MPC with privileged state and dynamics information, randomized over dynamics parameters and spatial disturbances. We then distill this expert MPC into an end-to-end visuomotor policy with a lightweight neural architecture, called SV-Net. SV-Net processes color image, optical flow and IMU data streams into low-level body rate and thrust commands at 20Hz onboard a drone. Crucially, SV-Net includes a Rapid Motor Adaptation (RMA) module that adapts at runtime to variations in drone dynamics. In a campaign of 105 hardware experiments, we show SOUS VIDE policies to be robust to 30% mass variations, 40 m/s wind gusts, 60% changes in ambient brightness, shifting or removing objects from the scene, and people moving aggressively through the drone’s visual field. Code, data, and experiment videos can be found on our project page: https://stanfordmsl.github.io/SousVide/.
@@ I INTRODUCTION
Learned visuomotor policies offer a compelling alternative to traditional drone navigation stacks by unifying perception and control into a streamlined framework. Unfortunately, training policies with human-like agility and collision avoidance requires a large corpus of visual and state data, making behavior cloning from human pilot demonstrations impractical. Simulation provides a promising alternative, but the sim-to-real gap has remained a persistent obstacle to real-world deployment. When overcome, these policies have shown highly agile, superhuman performance [1, 2, 3] However, such successes are limited to racing tracks designed for easy simulation reconstruction or hand-crafted simulation environments built from graphics asset libraries. This raises the question: how can we train an agile visuomotor policy to navigate cluttered real-world environments without relying on human input?
Figure 1: SOUS VIDE: Our simulator, FiGS, combines a GSplat scene model with a lightweight drone dynamics model. Within FiGS, an MPC policy generates demonstrations that track a desired trajectory across randomized drone dynamics and spatial disturbances. These are used to train a visuomotor policy, SV-Net, that we deploy zero-shot in the real-world.
@@
We address this challenge with FiGS (Flying in Gaussian Splats), a photorealistic drone simulator combining a Gaussian Splatting (GSplat) [4] scene model with a lightweight 9-dimensional drone dynamics model using thrust and body rate inputs—equivalent to the Acro mode used by expert human pilots. FiGS reconstructs scenes from handheld video captures with standard tools [5] and generates realistic image sequences and state estimation data for a drone at over 100 fps, all within an hour of acquiring video data on a standard GPU. This contrasts with the current practice of approximating each real scene with a handcrafted simulation instance [6, 7, 8, 9, 10, 11] that can take weeks of laborious sim-to-real transfer to perfect.
Building on FiGS, we develop a complete behavior cloning policy training pipeline and policy architecture for drone navigation, shown in Fig. 1. We call the combined simulator, training pipeline, and policy SOUS VIDE (Scene Optimized Understanding via Synthesized Visual Inertial Data from Experts). With SOUS VIDE, we can train a visuomotor policy to robustly navigate a drone in a real environment zero-shot—the policy has no access to real flight data at training time. Indeed, we can train policies to navigate drones through environments that neither the drone nor the users have previously visited, using remotely captured video or publicly available recordings.
@@
Specifically, we use FiGS to generate 100k-300k image/state-action pairs from an expert MPC policy following a desired nominal trajectory. The MPC has access to the ground truth state in the simulator and is therefore able to demonstrate high-quality, collision-free trajectories. To obtain stable and robust flight, we randomize dynamics parameters and spatial perturbations and record the MPC’s response. We use these expert demonstrations to train a student policy without privileged information (e.g. no lateral position data).
@@
For the student policy, we introduce a new architecture, called SV-Net, that is rich enough to ingest image and observable state history, while being efficient enough to run at 20Hz on board the drone. The policy ingests images using a SqueezeNet [12], re-trained on our data, and outputs a feature vector that is fused with observable data across several small MLPs to produce control commands. Within these, we also implement a form of the Rapid Motor Adaptation (RMA) concept from [13], which takes a history of observable data from a sliding window and produces a latent code that captures evolving flight dynamics in real time. We find this RMA module is crucial for robust real-world flight, adapting online to variations such as battery drain, rotor downwash effects and wind gusts.
To summarize, we introduce SOUS VIDE, which incorporates the following contributions:
@@
    Flying in Gaussian Splats (FiGS): A simulator coupling GSplat scene models with lightweight drone dynamics for photorealistic visual-inertial flight data.
@@
    Expert Demonstration Data: We use an MPC expert to synthesize behavior cloning data in FiGS with randomized dynamics parameters and positional disturbances.
@@
    SV-Net: A robust, efficient policy combining image and observable data to infer thrust and body rates, with RMA enabling online adaptation to changing flight conditions.
@@ II Related Work
Drone Polices Trained with GSplats: We are inspired by the recent works [14, 15] that train drone policies with visual data from a GSplat or Neural Radiance Field (NeRF) scene model. Most similar to our method, [14] uses a NeRF and an MPC expert to train a policy for trajectory following, but is not zero-shot (it requires real flight data) and it does not navigate through cluttered scenes. A zero-shot policy trained in a GSplat is presented in [15], but it is limited to tracking pre-placed visual markers with high-level commands that do not react to the obstacles in the scene. Some works have used GSplats for manipulation [16, 17], and locomotion [18] policies, but they are not applied to drone flight. To the best of our knowledge, SOUS VIDE is the first method to use GSplats to produce zero-shot drone navigation policies.
@@
High-performance simulation-trained policies have been demonstrated for drone racing [1, 2, 3], marking an impressive technological achievement. However, these methods blend real and simulation flight data, physics and learning-based models, and hand-engineered visual features cued into racing gates. In contrast, our method can train a policy using a casual video capture of any environment and can transfer zero-shot to the real-world scene without hand-tuned features.
Rapid Motor Adaptation (RMA): We borrow the concept of Rapid Motor Adaptation, originally developed for quadruped locomotion policies in [13]. RMA can be viewed as a pre-trained alternative to online parameter estimation [24]. RMA trains an encoder from sensing history to a latent vector that captures runtime operating conditions (e.g., terrain for a quadruped, or flight dynamics for a drone). RMA has been adapted for drones in [25, 26]. These policies achieve stable flight with impressive robustness but are not designed for visual navigation. Our lightweight RMA implementation is crucial for real-world robustness, addressing variations in both modeled and unmodeled drone dynamics.
Generalist Collision Avoidance Policies: Some existing works train policies to steer a drone through environments not seen at training time, often focusing on a particular scene domain like forests [27], office buildings [28], or urban roadways [29]. Such policies have been trained both with RL [28] and with imitation learning [27, 29], using both simulated [28, 23] and real-world [27, 30, 31, 32] data. Recent examples strive toward policies that can operate across different robot embodiments [31, 32]. While impressive for their generality, they often treat the drone as a pseudo-static ground robot by using a finely tuned onboard Visual-Inertial Odometry (VIO) stack to constrain the dynamics to planar velocities and yaw. This fails to exploit the drone’s full agility when navigating cluttered indoor environments with complex 3D trajectories. In contrast, SOUS VIDE directly commands thrust and body rates, mirroring the capabilities of expert human pilots.
@@ III Flying in Gaussian Splats (FiGS)
Here we describe the main components of our lightweight GSplat-based flight simulator, FiGS. We first collect a sequence of environment training images and train a GSplat model. We then forward integrate a simplified 9-dimensional drone dynamics model and render images from the GSplat as viewed through the onboard camera.
@@
Gaussian Splats: 3D Gaussian Splats [4] represent a 3D scene with a large collection of 3D Gaussians (up to millions), parameterized by mean $\mu\in\mathbf{R}^{3}$, covariance matrix $\Sigma\in\mathbf{S}_{++}$, opacity $\alpha\in\left[0,1\right]$, and spherical harmonics for view-dependent color, to approximate the geometry and appearance of real-world scenes. GSplats are coupled with a high-speed projection-based differentiable image rendering method called rasterization. They are trained from a sparse set of RGB images by backpropagating through the rasterization rendering to minimize photometric error between rendered images and training images. This approach achieves photorealistic reconstructions and renders full-resolution images at over 100 fps on a standard desktop GPU.
@@
In this work, we generate GSplats from short video recordings (1-2 minutes) of scene walk-throughs with a handheld camera. From the video we obtain a set of training images $(\mathcal{I}^{gt}_{0},\ldots,\mathcal{I}^{gt}_{N})$ and train the GSplat with the standard open-source tool Nerfstudio [5, 33]. The trained GSplat, $\mathcal{I}=\text{GS}_{\phi}(\bm{p},\bf{q})$, parametrized by $\phi$, can render photorealistic images $\mathcal{I}$ from a virtual camera placed at any pose (translation and quaternion rotation $(\mathbf{p},\mathbf{q})$) in the scene within the area covered by the training images. We obtain metric scale and align the GSplat frame to a known global frame in the scene by placing an ArUco tag marker on the ground during training.
@@
Drone Dynamics Model: Our model operates in the world, body, and camera frames ($\mathcal{W}$, $\mathcal{B}$,$\mathcal{C}$) and uses a simplified 9-dimensional semi-kinematic drone model. The states $\bm{x}=\begin{bmatrix}\bm{p}_{\mathcal{W}}&\bm{v}_{\mathcal{W}}&\bm{q}_{\mathcal{BW}}\end{bmatrix}^{T}$, represent translation ($\bm{p}_{\mathcal{W}}$), velocity ($\bm{v}_{\mathcal{W}}$) and quaternion orientation ( $\bm{q}_{\mathcal{BW}}$). The control inputs $\bm{u}=\begin{bmatrix}f_{th}&\bm{\omega}_{\mathcal{B}}\end{bmatrix}^{T}$ include normalized thrust ($f_{th}$) and desired angular velocities ($\bm{\omega}_{\mathcal{B}}$), forming the dynamics model
$\begin{split}\bm{\dot{p}}_{\mathcal{W}}&=\bm{v}_{\mathcal{W}},\\[-4.30554pt] \bm{\dot{v}}_{\mathcal{W}}&=-g\bm{z}_{\mathcal{W}}+k_{th}\frac{\hat{f}_{th}}{m_{dr}}\bm{z}_{\mathcal{B}}\\[-4.30554pt] \bm{\dot{q}}_{\mathcal{BW}}&=\frac{1}{2}\bm{W}(\bm{\omega}_{\mathcal{B}})\bm{q}_{\mathcal{BW}},\end{split},$ (1)
@@
where $g$ is gravitational acceleration, $\bm{W}(\bm{\omega}_{\mathcal{B}})$ is the quaternion multiplication matrix, and $\bm{z}_{\mathcal{W}}$ and $\bm{z}_{\mathcal{B}}$ are the world frame and body frame z-axis respectively. The model parameters $\bm{\theta}=\begin{bmatrix}k_{th}&m_{dr}\end{bmatrix}$ represent the thrust coefficient and drone mass.
@@
Body rate commands are the standard low-level input for most flight controllers [21, 1, 2, 3], providing robust tracking through high-rate gyroscope feedback. This choice also enhances platform agnosticism in a cost-effective manner and is widely favored by expert human pilots. Moreover, for our specific use case, it offers the significant advantage of omitting the rotational acceleration equations (Euler’s equations) in (1).
We forward integrate these equations of motion using ACADOS [34], a highly efficient trajectory optimizer that provides direct access to its integrator, to obtain state and control trajectories $(\bm{u}_{0:N-1},\bm{x}_{0:N})$. Applying a body-camera transform, we can render an image sequence from the GSplat $\mathcal{I}_{0:N}$ as seen by the onboard camera to yield data for policy training. This data can be used in a Reinforcement Learning (RL) or Imitation Learning (IL) framework, and can supervise training of either state-feedback or image-feedback policies. For SOUS VIDE, we use an IL framework with an MPC expert from ACADOS, which we will describe next.
@@ IV MPC Expert and Data Synthesis
Figure 2: Dynamic rollout of 50 data samples. At each time step, the update equation $f_{d}$ simulates the solution from the MPC expert, while the transform $T_{\mathcal{C}}^{\mathcal{B}}$ is used to extract the corresponding camera image $\mathcal{I}$ from the GSplat.
We begin by generating a desired trajectory through solving the minimum snap trajectory problem outlined in [35], which takes a sequence of position-yaw waypoints and outputs a dynamically feasible spline, from which a state trajectory and open-loop control trajectory are obtained through the differential flatness map. This waypoint-based approach allows us to interface with various high-level planners commonly used in drone trajectory planning. By sampling at our policy’s control frequency, $\nu_{ctl}$, we produce a $N_{des}$ step trajectory $\mathbf{X}^{\text{des}}=\{\bm{x}^{des}_{0},\cdots\bm{x}^{des}_{N_{des}}\}$ with control inputs parametrized by the drone dynamics parameters $\bm{\theta}$, $\mathbf{U}^{\text{des}}=\{\bm{u}^{des}_{0},\cdots\bm{u}^{des}_{N_{des}-1}\}$.
A large demonstration dataset can be generated using any trajectory optimization technique with dynamics rollouts, such as the one-shot sampling methods in [36] or by iteratively running DAgger [37] in simulation. We opt for the simplest approach, domain randomization, as described in Algo. 1 and illustrated in Fig. 2. This leverages the strength of FiGS in quickly producing large volumes of photorealistic image data while leaving the door open to more sophisticated techniques.
We define $(\bm{\theta}_{\text{min}},\bm{\theta}_{\text{max}},\Delta\bm{x})$ as the bounds on the uniform distributions on the drone parameters and sample states, $N_{\text{smp}}$ is the step-wise sample count and $\mathcal{GS}$ is the Gaussian Splat. We observe that these bounds can be liberally set. GenerateDynamics instantiates (1) with a given set of drone parameters ($f_{d}$ in Fig. 2), GenerateImage extracts the camera pose given the drone state and camera transform ($T_{\mathcal{C}}^{\mathcal{B}}$ in Fig. 2) before rendering the corresponding image from the Gaussian Splat ($\mathcal{I}^{k}$ in Fig. 2). TMPC is our expert policy that uses MPC to track $X^{\text{des}}$ and $U^{\text{des}}$, starting from the nearest position to $\bm{x}_{s}$ for $t_{\text{smp}}$ seconds. We optimize
$\begin{split}&\min_{\bm{u}}\sum_{k=0}^{N-1}(\delta\bm{x}_{k}^{T}Q_{k}\delta\bm{x}_{k}+\delta\bm{u}_{k}^{T}R_{k}\delta\bm{u}_{k})+\delta\bm{x}_{N}^{T}Q_{N}\delta\bm{x}_{N}\\ &\text{s.t.}\quad\bm{x}_{k+1}=f_{d}(\bm{x}_{k},\bm{u}_{k}),\quad g_{c}(\bm{u}_{k})\leq\bm{0}\end{split}.$ (2)
in an $N$-step receding horizon manner, subject to dynamics constraint $f_{d}$ and the control limits constraint $g_{c}$. The stage-wise weights, $(Q_{k},R_{k})$, and the terminal weights, $Q_{N}$, are applied to the difference between the rollout trajectory and the desired trajectory, defined $\delta\bm{x}_{k}=(\bm{x}_{k}-\bm{x}^{des}_{k})$, and similarly for $\delta\bm{u}_{k}$. We use ACADOS to implement the MPC using it, though alternative solvers could be employed.
@@
1:$\mathbf{X}^{\text{des}}$,$\mathbf{U}^{\text{des}}$,$\bm{\theta}_{\text{min}}$,$\bm{\theta}_{\text{max}}$,$\Delta\bm{x}$,$N_{\text{des}}$,$N_{\text{smp}}$,$t_{\text{smp}}$,$T_{C}^{B}$,$\mathcal{GS}$
@@
3:for $i=0$ to $N_{\text{des}}-1$ do
4:     for $j=0$ to $N_{\text{smp}}$ do
5:         $\bm{\theta}_{s}\!\sim\!U(\bm{\theta}_{\text{min}},\bm{\theta}_{\text{max}})$, $\bm{x}_{s}\!\sim\!(\bm{x}^{des}_{i}-\Delta\bm{x},\bm{x}^{des}_{i}+\Delta\bm{x})$
@@
7:         $\mathbf{X}^{\text{smp}},\mathbf{U}^{\text{smp}}=\texttt{TMPC}(\bm{x}_{s},f_{d},t_{\text{smp}},\mathbf{X}^{\text{des}},\mathbf{U}^{\text{des}})$
8:         $\mathbf{I}^{\text{smp}}=\texttt{GenerateImage}(\mathbf{X}^{\text{smp}},T_{C}^{B},\mathcal{GS})$
9:         $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathbf{X}^{\text{smp}},\mathbf{U}^{\text{smp}},\mathbf{I}^{\text{smp}})\}$

### v2

@@ Abstract
We propose a new simulator, training approach, and policy architecture, collectively called SOUS VIDE, for end-to-end visual drone navigation. Our trained policies exhibit zero-shot sim-to-real transfer with robust real-world performance using only onboard perception and computation. Our simulator, called FiGS, couples a computationally simple drone dynamics model with a high visual fidelity Gaussian Splatting scene reconstruction. FiGS can quickly simulate drone flights producing photorealistic images at up to 130 fps. We use FiGS to collect 100k-300k image/state-action pairs from an expert MPC with privileged state and dynamics information, randomized over dynamics parameters and spatial disturbances. We then distill this expert MPC into an end-to-end visuomotor policy with a lightweight neural architecture, called SV-Net. SV-Net processes color image, optical flow and IMU data streams into low-level thrust and body rate commands at 20 Hz onboard a drone. Crucially, SV-Net includes a learned module for low-level control that adapts at runtime to variations in drone dynamics. In a campaign of 105 hardware experiments, we show SOUS VIDE policies to be robust to 30% mass variations, 40 m/s wind gusts, 60% changes in ambient brightness, shifting or removing objects from the scene, and people moving aggressively through the drone’s visual field. Code, data, and experiment videos can be found on our project page: https://stanfordmsl.github.io/SousVide/.
@@ I INTRODUCTION
Learned visuomotor policies offer a compelling alternative to traditional drone navigation stacks by unifying perception and control into a streamlined framework. Unfortunately, training policies with human-like agility and collision avoidance requires a large corpus of visual and state data, making behavior cloning from human pilot demonstrations impractical. Simulation provides a promising alternative, but the sim-to-real gap has remained a persistent obstacle to real-world deployment. Recent work has demonstrated that in controlled environments and with carefully built digital twins in simulation, learned policies can achieve highly agile, superhuman performance [1, 2, 3]. However, this raises the question: can we train a visuomotor policy to navigate unstructured real-world environments with minimal human curation?
Figure 1: SOUS VIDE overview: We train our FiGS simulator from a hand held camera. We use FiGS to generate flight demonstrations (image/state-action pairs) from an MPC expert with privileged information randomized over dynamics parameters and positional disturbances. We use this data to train our policy, SV-Net, which operates solely with onboard observations.
@@
We address this challenge with FiGS (Flying in Gaussian Splats), a photorealistic drone simulator combining a Gaussian Splat (GSplat) [4] scene model with a lightweight 10-dimensional drone dynamics model using thrust and body rate inputs—equivalent to the Acro mode used by expert human pilots. FiGS reconstructs scenes from video captures, like publicly available footage or smartphone recordings, processed with standard tools [5] and generates realistic image sequences and state estimation data for a drone at up to 130 fps on a standard GPU, all within an hour of acquiring video data. This contrasts with the current practice of approximating each real scene with a handcrafted simulation instance [6, 7, 8, 9, 10, 11] that can take weeks of laborious sim-to-real transfer to perfect.
@@
Building on FiGS, we develop SOUS VIDE [^2] The name is inspired by the sous vide cooking technique that evenly cooks food in a vacuum-sealed bag under carefully controlled conditions. (Scene Optimized Understanding via Synthesized Visual Inertial Data from Experts), a behavior cloning pipeline that produces a robust drone navigation policy capable of zero-shot sim-to-real transfer—trained entirely within simulation without real-world demonstrations or fine-tuning (Fig. 1). Specifically, we use FiGS to generate 100k-300k image/state-action pairs from an expert MPC policy following a desired nominal trajectory within a GSplat. The MPC has access to the ground truth state in the simulator and is therefore able to demonstrate high-quality, collision-free trajectories. To obtain stable and robust flight, we randomize dynamics parameters and spatial perturbations and record the MPC’s response. We use these expert demonstrations to train a student policy without privileged information (no lateral position data).
@@
The learned policy produced by SOUS VIDE, SV-Net, is a novel architecture designed to process both images and observable state history while remaining efficient enough to run onboard the drone. The policy ingests images using a SqueezeNet [12], re-trained on our data, and outputs a feature vector that is fused with observable data across several small Multi-Layer Perceptrons (MLPs) to produce low-level body-rate commands. Within these MLPs, we also implement a form of the Rapid Motor Adaptation (RMA) concept from [13], which takes a history of observable data from a sliding window and produces a latent code that captures evolving flight dynamics in real time. We find this RMA module is crucial for robust flight, adapting online to variations such as battery drain, rotor downwash effects, and wind gusts.
To summarize, we make the following contributions:
@@
    Flying in Gaussian Splats (FiGS): A simulator coupling GSplat scene models with drone dynamics for efficient and photorealistic visual-inertial data generation.
@@
    Scalable Visuomotor Policy Generation: We use FiGS to generate large synthetic datasets to train visuomotor policies that transfer zero-shot to real-world flight.
@@
    SV-Net: An onboard policy that fuses image and observable states to infer thrust and body rates while continuously adapting to changing flight conditions.
@@ II Related Work
Drone Polices Trained with GSplats: Learned representations like GSplats have proven effective in training visuomotor policies across many domains, from manipulation [14, 15] to bipedal locomotion [16] to aerial robotics [17, 18]. Closest to our approach, [17] uses a learned representation and an MPC expert to train a trajectory-following policy but requires an initial sample of real-world expert flight demonstrations to be collected via motion capture. Additionally, its $45^{\circ}$ downward-facing camera focuses on ground features, missing the spatial information of a forward view needed for obstacle avoidance, as seen in drone racing [2]. Meanwhile, [18], treats the GSplat reconstruction as a background, relying instead on colored spheres as visual markers injected into both the simulation and real-world scenes as the basis for decision making. Moreover, it relies on velocity commands that encode only high-level approach and turn behaviors, delegating low-level control to a manufacturer-supplied autonomy stack. To the best of our knowledge, SOUS VIDE is the first method to leverage GSplats for generating low-level drone navigation policies for unstructured environments without assistive infrastructure or real-world expert flight data.
@@
High-performance simulation-trained policies have been demonstrated for drone racing [2, 3], marking an impressive technological achievement. However, these methods blend real and simulation flight data, physics and learning-based models, and hand-engineered visual features cued into racing gates. In contrast, our method can train a policy using video clips of the scene and can transfer zero-shot to the real-world with only minimal tuning of easily measurable parameters.
Rapid Motor Adaptation: RMA, a technique originally developed for quadruped locomotion policies in [13], can be viewed as a pre-trained alternative to online parameter estimation [24] where an encoder is trained to take in a sensing history to produce a latent vector that captures runtime operating conditions (e.g., terrain for a quadruped, or flight dynamics for a drone). RMA has been adapted for drones in [25, 26] where they have been show to achieve stable flight with impressive robustness. However, they are not designed for visual navigation. Our lightweight RMA implementation is crucial for real-world robustness, addressing variations in both modeled and unmodeled drone dynamics.
Generalist Collision Avoidance Policies: Some existing works train policies to steer a drone through environments not seen at training time, often focusing on a particular scene domain like forests [27], office buildings [28], or urban roadways [29]. Such policies have been trained both with Reinforcement Learning (RL) [28] and with Behavior Cloning (BC) [27, 29], using both simulated [28, 23] and real-world [27, 30, 31, 32] data. Recent examples strive toward policies that can operate across different robot embodiments [31, 32]. While impressive for their generality, they often treat the drone as a pseudo-static ground robot by using a finely tuned onboard Visual-Inertial Odometry (VIO) stack to constrain the dynamics to planar velocities and yaw. This fails to exploit the drone’s full agility when navigating cluttered indoor environments with complex 3D trajectories. In contrast, SOUS VIDE directly commands thrust and body rates, mirroring the capabilities of expert human pilots.
@@ III Flying in Gaussian Splats (FiGS)
FiGS, our lightweight GSplat-based flight simulator, consists of a GSplat model trained from video captures of the scene, within which a drone is simulated using a simplified 10-dimensional drone dynamics model.
@@
Gaussian Splats: 3D Gaussian Splatting [4] is a learned representation approach that approximates the geometry and appearance of real-world scenes using a large collection of Gaussians—potentially millions—each parameterized by its position, covariance, color, and opacity. They leverage high-speed, projection-based differentiable rasterization and are trained from sparse RGB images by backpropagating through the rasterization to minimize photometric error. This approach enables photorealistic reconstructions and full-resolution renders at over 100 fps on a standard desktop GPU.
@@
In this work, we generate GSplats from short video recordings (2-3 minutes) of scene walk-throughs with a handheld camera. From the video we extract a set of training images and use the open-source tool Nerfstudio [5, 33] to train the GSplat model. The resulting model $\mathcal{GS}_{\phi}$, where $\phi$ are its parameters, can render photorealistic images from a virtual camera placed at any pose within the region covered by the training images. Given a camera pose $(\bm{p},\bm{q})$, where $\bm{p}$ represents the position and $\bm{q}$ the orientation in quaternion form, the rendered image is given by $\bm{I}=\mathcal{GS}_{\phi}(\bm{p},\bf{q})$. To obtain metric scale and align the GSplat frame to a known global frame in the scene, we start the video recording with an ArUco tag marker in frame.
@@
Drone Dynamics Model: Our model operates in the world, body, and camera frames ($\mathcal{W}$, $\mathcal{B}$, $\mathcal{C}$) and uses a 10-dimensional semi-kinematic state vector, $\bm{x}=\begin{bmatrix}\bm{p}_{\mathcal{W}},\bm{v}_{\mathcal{W}},\bm{q}_{\mathcal{BW}}\end{bmatrix}^{T}$, representing position $\bm{p}_{\mathcal{W}}=(p_{x},p_{y},p_{z})$, velocity $\bm{v}_{\mathcal{W}}=(v_{x},v_{y},v_{z})$, and orientation $\bm{q}_{\mathcal{BW}}=(q_{x},q_{y},q_{z},q_{w})$. The control inputs, $\bm{u}=\begin{bmatrix}f_{th},\bm{\omega}_{\mathcal{B}}\end{bmatrix}^{T}$, include normalized thrust $f_{th}$ and angular velocity $\bm{\omega}_{\mathcal{B}}=(\omega_{x},\omega_{y},\omega_{z})$. This produces model dynamics
$\begin{split}\bm{\dot{p}}_{\mathcal{W}}&=\bm{v}_{\mathcal{W}},\\[-4.30554pt] \bm{\dot{v}}_{\mathcal{W}}&=g\bm{z}_{\mathcal{W}}-k_{th}\frac{f_{th}}{m_{dr}}\bm{z}_{\mathcal{B}}\\[-4.30554pt] \bm{\dot{q}}_{\mathcal{BW}}&=\frac{1}{2}\bm{W}(\bm{\omega}_{\mathcal{B}})\bm{q}_{\mathcal{BW}},\end{split},$ (1)
@@
where $g$ is gravitational acceleration, $\bm{W}(\bm{\omega}_{\mathcal{B}})$ is the quaternion multiplication matrix, and $\bm{z}_{\mathcal{W}}$, $\bm{z}_{\mathcal{B}}$ are the z-axis unit vectors of the world and body frames. The thrust coefficient and mass, $(k_{th},m_{dr})$, are stored in the drone parameter vector $\bm{\theta}$.
@@
Thrust and body rate commands are the standard low-level input for most flight controllers [21, 1, 2, 3], providing robust tracking through high-rate gyroscope feedback. This choice also enhances platform agnosticism in a cost-effective manner and is widely favored by expert human pilots. Moreover, for our use case, it offers the significant advantage of omitting the rotational acceleration equations (Euler’s equations) in (1).
[…]
@@ IV MPC Expert and Data Synthesis
Figure 2: Dynamic rollout of 50 data samples. At each time step, the update function $f_{d}$ simulates the solution from the MPC expert, while the transform $T_{\mathcal{C}}^{\mathcal{B}}$ is used to extract the corresponding camera image $I$ from the GSplat.
Many drone navigation frameworks are designed around a desired trajectory, whether to encode complex paths [35, 36, 17], race courses over a sequence of gates [2, 3], or even optimal approaches for perching [37]. Given the complexity and variety of motion planning objectives, this abstraction facilitates a decoupled approach where high-level goals can be achieved by a higher-level task planner that generates a desired trajectory for a low-level navigation policy to execute. For instance, if the goal is obstacle avoidance, one could use the already existing GSplat to generate collision-free waypoints [38] that could then be turned into a desired trajectory.
In this work we are interested in the ability to navigate tight spaces and so we handpick a sequence of waypoints that intentionally guides the drone near or through obstacles. From these we use [35] to compute a dynamically feasible spline which we then sample at our desired control frequency $\nu_{\text{ctl}}$ to extract an $N_{d}$-step state and input desired trajectory $(\mathbf{X}^{d},\mathbf{U}^{d})$, parametrized by $\bm{\theta}$. We can then apply one of a variety of trajectory optimization techniques, such as the one-shot sampling methods in [36] or even an iterative form of DAgger [39], to guide a drone towards the desired trajectory in simulation. We opt for the simplest approach, domain randomization, as described in Algo. 1 and illustrated in Fig. 2. This leverages the strength of FiGS in quickly producing large volumes of photorealistic image data while leaving the door open to more sophisticated techniques.
[…]
$\begin{split}&\min_{\bm{u}}\sum_{k=0}^{N-1}(\delta\bm{x}_{k}^{T}Q_{k}\delta\bm{x}_{k}+\delta\bm{u}_{k}^{T}R_{k}\delta\bm{u}_{k})+\delta\bm{x}_{N}^{T}Q_{N}\delta\bm{x}_{N}\\ &\text{s.t.}\quad\bm{x}^{s}_{k+1}=f_{d}(\bm{x}^{s}_{k},\bm{u}^{s}_{k}),\quad g_{c}(\bm{u}^{s}_{k})\leq\bm{0}\end{split}.\vspace{-1mm}$ (2)
[…]
@@
1:$\mathcal{GS}_{\phi}$, $\mathbf{X}^{d}$, $\mathbf{U}^{d}$, $N_{d}$, $\bm{\theta}_{\text{min}}$, $\bm{\theta}_{\text{max}}$, $\Delta\bm{x}$, $N_{s}$, $t_{s}$, $T_{\mathcal{C}}^{\mathcal{B}}$
@@
3:for $i=0$ to $N_{d}$ do
4:     for $j=0$ to $N_{s}$ do
5:         $\bm{\theta}_{s}\!\sim\!U(\bm{\theta}_{\text{min}},\bm{\theta}_{\text{max}})$, $\bm{x}^{s}_{0}\!\sim\!(\bm{x}^{d}_{i}-\Delta\bm{x},\bm{x}^{d}_{i}+\Delta\bm{x})$
@@
7:         $\mathbf{X}^{s},\mathbf{U}^{s}=\texttt{MPC}(\bm{x}^{s}_{0},f_{d},t_{s},\mathbf{X}^{d},\mathbf{U}^{d})$
[…]